OpenStudy (anonymous):

Help please! I only have a few more problems b4 i can got to bed! :D lol thank you <3 What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)? y = 2(x + 2) y = 2(x – 2) y = –2(x – 2) y = –2(x + 2)

OpenStudy (anonymous):

its (y-0)=(8-0)/(2+2) (x+2) or y=2(x+2)

OpenStudy (anonymous):

(y-0)=(8-0)/(2+2) (x+2)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

first option

OpenStudy (anonymous):

Ya'll are giving me 2 different answers :/ lol

OpenStudy (anonymous):

sorry i meant the first one but a d came out sorry

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Last question! y varies directly as the square of x. When x = 1, y = 8. Find x when y = 128. x = ±4 x = ±6 x = ±8 x = ±16

OpenStudy (anonymous):

Ok! :D

OpenStudy (nurali):

y = 8x² 128 = 8x² 16 = x² x = +-4

OpenStudy (nurali):

if correct please click best response.

OpenStudy (anonymous):

Hold up.. I lied..I got one more.. lol. Which inequality is graphed?Graph of an absolute value inequality with the vertex on three-halves comma one and shaded below. A dotted line crosses through points (1, 0) and (2, 0). y > – |2x + 3| – 1 y < |x + 3| + 1 y < – |2x – 3| + 1 y < |2x – 3| – 1

OpenStudy (anonymous):

ok :/

OpenStudy (anonymous):

so vertex (3/2,1)

OpenStudy (anonymous):

Which one? lol

OpenStudy (anonymous):

i am asking about the vertex? is that what you meant lol

OpenStudy (anonymous):

oooooh. haha. Uhm surreee :P lol

OpenStudy (anonymous):

lol alright that point is what you are goin to need to figure it out. if it is 3/2, 1 then it should be i believe the 3rd one

OpenStudy (anonymous):

Thank youuu <3

OpenStudy (anonymous):

mhmm:)

OpenStudy (aravindg):

@magenxoxo still cant message u or fan u !! :(

OpenStudy (aravindg):

tday i gt my modem fixed up ,tell me when u r here :)